OpenStudy (anonymous):

2 2 3 _ - __ + ___ x x-1 (x-1)^2 HELP!!!! :(

Directrix (directrix):

I am not sure what the problem is. Would you click on the blue DRAW tab below and write the problem? Thanks.

OpenStudy (anonymous):

\[\frac{ 2 }{ x }-\frac{ 2 }{ x-1 }+\frac{ 3 }{ (x-1)^2 }\] is that what you were trying to enter?

OpenStudy (anonymous):

YES!!! :)

OpenStudy (anonymous):

HELP!!!

Directrix (directrix):

What are the instructions? Write as a single fraction, or maybe simplify.

OpenStudy (anonymous):

Add or subtract as indicated and simplify the result.

Directrix (directrix):

We will need a common denominator for x, x-1, and (x-1)².

OpenStudy (anonymous):

Would it be x(x-1)^2?

Directrix (directrix):

Yes. The task now is to rewrite each of the fractions in the problem with that denominator.

Directrix (directrix):

What is 2/x equal to when written with a denominator of x(x-1)^2 ?

Directrix (directrix):

Go ahead.

OpenStudy (anonymous):

2(x-1)^2

OpenStudy (anonymous):

So that would mean that it would be 2x^2-4x+2 over the LCD?

Directrix (directrix):